Julie has three children whose ages are consecutive odd integers.
If x represents the youngest child’s age, which expression represents
the sum of her children’s ages?

Answer by JulietG(1812)

"Which expression" gives the opinion that there are choices, yet you did not list any.
.
Let A, B, and C be the children's ages.
The children are:
A (oldest child)
B = A-2 (the numbers are consecutive and odd, so 2 apart)
C = A-4
.
The sum is A + B + C
Substitute the values above.
A + (A-2) + (A-4)
Add
3A - 6 = sum of children's ages